OpenStudy (yellowfellow007):

A tractor is pulling two logs in tandem. The logs are fastened together by a chain. Each log weighs 1800.0 N, and the logs are moving with a constant velocity. The tractor exerts a force of 1440 N on the front log, and this same log is acted on by a force of friction of 941 N.

OpenStudy (yellowfellow007):

Determine the tension of the rope between the logs, determine the force of friction on the back logs

OpenStudy (yellowfellow007):

BTW, I calculated that mu= 0.4 . constant velocity=accleration=0. a=f/m 0=f/m. f=0. f 1440N-Ffrict=0. ffrict= mu*fnorm. 1440-3600N(mu)=0 mu=0.4